• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Wouter Deconinck

    Log entry time 08:00:32 on December13,2008

    Entry number 253190

    keyword=owl shift summary

    Summary owl shift 13-Dec-2008

    Shift crew:

    S.L.: Wouter Deconinck
    T.O.: Wouter Deconinck
    3rd: Salvatore Frullani


    Spectrometer:

    LHRS momentum = 2.35 GeV/c
    LHRS angle = 16 deg
    BigBite sieve = OUT
    Target collimator = IN
    lambda/2 plate = IN
    Field direction = trans
    Target = 3He pol


    --00:00 Shift starts, run 4434 in progress

    RUN START STOP COULOMB
    4434 23:04 00:06 0.04006
    4435 00:07 01:11 0.04004
    4436 01:12 02:21 0.03979
    --02:22 HAPPEX run 30563 started
    4437 02:22 03:26 0.04005
    4438 03:37 04:30 0.04018
    --04:14 Oven temperature drops to 200 deg due to PID glitch
    4439 04:30 04:46 0.008161
    --04:44 BigBite L14.6 and RICH ch 16 trip during run 4439
    RICH does not recover
    --04:44 ROC3 locked up, not able to revive using software
    --05:10 DAQ expert advises going in the hall to power cycle ROC3
    --05:37 ROC3 power cycle does not solve problem
    --06:00 Problem with 5V power supply in the ROC3 -> down until 9am
    Parasitic access for target flip
    RICH experts unreachable

    --08:00 Shift ends, experiment down until 9am

    Accumulated 0.20828 C in 6 runs.